# From graph to Riesz continuity

### v2

@@ 1 Introduction
When all the operators $\mathcal{A}_{x}$ are contained in a fixed connected component of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}$, there is another kind of local obstruction which we now describe. Let $\lambda\in\mathbb{R}$ be chosen outside of the spectrum of $\mathcal{A}_{x_{0}}$, $x_{0}\in X$. Then the spectral projections $p_{x}=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}(\mathcal{A}_{x})$ (here $\operatorname{\mathbb{1}}_{S}$ denotes the characteristic function of a subset $S\subset\mathbb{R}$) give rise to the Hilbert bundles $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$, with fibers $\mathcal{H}^{+}_{x}=\operatorname{Ran}(p_{x})$ and $\mathcal{H}^{-}_{x}=\operatorname{Ker}(p_{x})$, over some neighborhood of $x_{0}$. A local adapted trivialization of $\mathcal{H}$ near $x_{0}$ exists if and only if both $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ are locally trivial near $x_{0}$. See Proposition 5.3.
@@
Suppose first that all the operators $\mathcal{A}_{x}$ are essentially negative. Since the norm and strong topologies coincide on the space of projections of a fixed finite rank, an obstruction to existence of local adapted trivializations near $x_{0}$ arises when the rank of $p_{x}$ is not locally constant near $x_{0}$. In other words, there are eigenvalues of $\mathcal{A}_{x}$ going to +$\infty$+ ∞ when $x\to x_{0}$. If this is not the case and the operators $\mathcal{A}_{x}$ are uniformly bounded from above in a neighborhood of each point $x_{0}\in X$, then $\mathcal{A}$ is Riesz continuous in every trivialization of $\mathcal{H}$. The situation with essentially positive operators is completely the same, but +$\infty$+ ∞ and bounds from above should be replaced by $\infty$- ∞ and bounds from below. See, e.g., Rellich’s example [Ka, Example V-4.14].
Suppose now that all $\mathcal{A}_{x}$ are contained in the third connected component of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}(H)$, ${{}^{r}\mathcal{R}}_{K}^{\star}(H)$, which consists of operators that are neither essentially positive nor essentially negative. Then both Hilbert bundles $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ have infinite rank. By a result of Dixmier and Douady, every separable Hilbert bundle of infinite rank over a paracompact base space of finite covering dimension is trivial. Recall that $X$ is said to have finite covering dimension if there is an $n\in\mathbb{N}$ such that every finite open covering of $X$ has a finite open refinement of order $\leqslant n$.
Let us say that a space has locally finite covering dimension if each point has a neighborhood of finite covering dimension. Using the aforementioned result of Dixmier and Douady, we prove a partial analogue of Theorem C for self-adjoint operators.
Theorem E.
@@
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a metrizable space $X$ of locally finite covering dimension. Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators $\mathcal{A}_{x}\in\mathcal{R}^{\star}_{K}(\mathcal{H}_{x})$, with compact resolvents, that are neither essentially positive nor essentially negative. Then $\mathcal{A}$ becomes Riesz continuous in some global trivialization of $\mathcal{H}$. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization.
@@
Since the group ${{}^{s}{\mathcal{U}}}(H)$ is contractible, Theorem E implies that the identity map ${{}^{r}\mathcal{R}}^{\star}_{K}\to{{}^{g}\mathcal{R}}^{\star}_{K}$ is a weak homotopy equivalence. The stronger result, of a homotopy equivalence of this map, is proven in the previous paper of the author by a different method. See [$P_{3}$], Theorems B and C.
@@
Theorem E does not hold without additional assumptions for general metrizable base spaces, even compact, as we demonstrate by Example 5.5.
Elliptic operators on closed manifolds.   Let us describe an important class of situations where Theorems C and D may be useful. Let $\mathcal{M}\to X$ be a smooth fibration whose fibers $\mathcal{M}_{x}$ are smooth closed manifolds, and let $\mathcal{E}$, $\mathcal{E}^{\prime}$ be smooth vector bundles over $\mathcal{M}$. Let $\mathcal{D}_{x}$ be a continuous (in an appropriate sense) family of elliptic pseudodifferential operators on $\mathcal{M}_{x}$, of positive order, acting from fibers of $\mathcal{E}$ to fibers of $\mathcal{E}^{\prime}$. A local trivialization of $\mathcal{M}$ turns $\mathcal{D}$ to a family of elliptic operators acting between fibers of fixed vector bundles $E$ and $E^{\prime}$ over a fixed closed manifold $M$. The bounded transform takes this new family to a continuous family of zeroth order elliptic operators, which is norm continuous as a family of bounded operators $L^{2}(M;E)\to L^{2}(M;E^{\prime})$. Hence the family of unbounded operators $\mathcal{D}_{x}$ is Riesz continuous in this local trivialization. However, the fibration $\mathcal{M}\to X$ is in general non-trivial, and thus no such “geometric” global trivialization exists in the general case. The structure group of the corresponding Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, where $\mathcal{H}_{x}=L^{2}(\mathcal{M}_{x};\mathcal{E}_{x})$ and $\mathcal{H}^{\prime}_{x}=L^{2}(\mathcal{M}_{x};\mathcal{E}^{\prime}_{x})$, has the topology coarser than the norm topology, so Kuiper’s theorem cannot be used in order to trivialize $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Dixmier-Douady theorem, on the other hand, can be used, but trivializations it provides do not take $\mathcal{D}$ to a Riesz continuous family of operators.
Theorem C closes this gap, showing that such a family $\mathcal{D}$ can be made Riesz continuous by appropriate global trivializations of the Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, though at the cost of losing the differential and geometric structure. Theorem D solves the same problem for families of self-adjoint elliptic operators acting on fibers of $\mathcal{E}$.
@@
The author arrived at Theorem D in an attempt to understand a step in the proof of a result of Melrose and Piazza. See [$MP_{1}$], Proposition 1. Melrose and Piazza deal with a family of self-adjoint differential operators of first order over a smooth fibration $\mathcal{M}\to X$, as above, and show that if the $K^{1}$-index of the family vanishes, then the family admits a spectral section. In their proof, they claim that the $L^{2}$ spaces of the fibers, $L^{2}(\mathcal{M}_{x};\mathcal{E}_{x})$, can be identified so that all operators act on a fixed Hilbert space $H$. Unfortunately, they did not explain how such an identification can be obtained. The author was not able to find a simple explanation of this step. Our Theorem D, as well as recent results [$I_{2}$, Theorems 4.5 or 4.6] or [$I_{3}$] of N. Ivanov, allows to overcome this difficulty. Similarly, Theorem C allows to pass to a fixed Hilbert space in the proof of [$MP_{2}$, Proposition 2], an odd analog of [$MP_{1}$, Proposition 1].
Elliptic operators on manifolds with boundary.   When the fibers $\mathcal{M}_{x}$ are compact manifolds with boundary, the situation becomes more complicated. In this case $\mathcal{D}_{x}$ is not even a closed operator usually. One needs to add an elliptic boundary condition $\mathcal{B}_{x}$ to $\mathcal{D}_{x}$, and then such a pair $(\mathcal{D}_{x},\mathcal{B}_{x})$ determines a regular operator with compact resolvents. Taking a local trivialization of $\mathcal{M}$ as above, one obtains a family of boundary value problems $(\mathcal{D}_{x},\mathcal{B}_{x})$ on a fixed manifold $M$. If both $\mathcal{D}_{x}$ and $\mathcal{B}_{x}$ depend continuously on $x$ in an appropriate sense, then the corresponding family of regular operators with compact resolvents is graph continuous. See [$P_{1}$, Appendix A.5] or [BZ, Corollary A.6.3]. Theorem C shows that such a family $\mathcal{D}$ can be made Riesz continuous by appropriate global trivializations of the corresponding Hilbert bundles.
Suppose now that we have a family of self-adjoint elliptic boundary value problems. As before, a local trivialization of $\mathcal{M}$ gives rise to a family of boundary value problems on a fixed manifold $M$. Unfortunately, there are no criteria of Riesz continuity for families of unbounded operators arising from boundary value problems, except for several special cases. Moreover, there are simple examples of such families that are not Riesz quasi-continuous, see e.g. Rellich’s example. Theorem E, however, requires only graph continuity of a family of unbounded operators, so it can be applied to families of elliptic self-adjoint boundary value problems, provided that they are neither essentially positive nor essentially negative and the space of parameters is good enough.
Acknowledgment.   I am grateful to N. V. Ivanov for attracting my attention to the work of Dixmier and Douady [DD] and for helpful remarks.
@@ 2 Preliminaries
Grassmanian.   The set $\operatorname{Gr}(H)$ of closed subspaces of $H$ is called the Grassmanian. An orthogonal projection may be identified with its range, which provides a bijection between $\operatorname{Gr}(H)$ and $\mathcal{P}(H)$. The Grassmanian is equipped with the topology (and even metric) induced by this bijection from the norm on the space $\mathcal{P}(H)$ of projections. We will use closed subspaces and projections by turn, depending on what is more convenient at the moment.
Canonical decomposition.   Recall a standard construction that will be used throughout the paper. The trivial Hilbert bundle over $\operatorname{Gr}(H)$ with the fiber $H$ is canonically decomposed into the direct sum
$\operatorname{Gr}(H)\times H=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$
of two Hilbert bundles, whose fibers over $L\in\operatorname{Gr}$ are $\mathcal{H}^{+}_{L}=L$ and $\mathcal{H}^{-}_{L}=L^{\bot}$, the orthogonal complement of $L$ in $H$. This decomposition is locally trivial in the following sense: for every $L_{0}\in\operatorname{Gr}(H)$ there is a continuous map $g\colon B\to\mathcal{U}(H)$, where $B$ is the open ball of radius $1$ around $L_{0}$, such that $L=g_{L}L_{0}$ for every $L\in B$. In terms of projections, this is written as $p=g_{p}p_{0}g_{p}^{*}$ for $\left\|p-p_{0}\right\|<1$, where $p_{0}$ is the projection onto $L_{0}$. See the proof of [Ka, Theorem I.6.32] for the construction of such a map $g$.
An important subspace of $\operatorname{Gr}(H)$ is
$\operatorname{Gr}^{\star}(H)=\left\{L\in\operatorname{Gr}(H)\,|\,\dim L=\infty=\operatorname{codim}L\right\}.$
Let $L_{0}\in\operatorname{Gr}^{\star}(H)$ be a fixed subspace. The map $\mathcal{U}(H)\to\operatorname{Gr}^{\star}(H)$, $u\mapsto uL_{0}$ is a locally trivial principal bundle with the contractible structure group $\mathcal{U}(L_{0})\times\mathcal{U}(L_{0}^{\bot})$ over a paracompact base space and therefore has a section
(2.1) $g\colon\operatorname{Gr}^{\star}(H)\to\mathcal{U}(H),\;\;\text{so that}\;\;L=g_{L}L_{0}\;\;\text{for all}\;\;L\in\operatorname{Gr}^{\star}(H).$
It follows that the restrictions of $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ to $\operatorname{Gr}^{\star}$ are trivial Hilbert bundles with the structure group $\mathcal{U}$. In addition, contractibility of the total space $\mathcal{U}(H)$ and the fiber $\mathcal{U}\times\mathcal{U}$ implies contractibility of the base space $\operatorname{Gr}^{\star}$. See [AS, proof of Lemma 3.6].
Bounded transform.   Let us recall the definition and the main properties of the bounded transform. See, e.g., [Sc], Theorem 7.5 for the detail.
@@
defines the inclusion of the set $\mathcal{R}(H,H^{\prime})$ of regular operators to the closed unit ball
$\mathcal{D}(H,H^{\prime})=\left\{a\in\mathcal{B}(H,H^{\prime})\,|\,\left\|a\right\|\leqslant 1\right\}.$
@@
The orthogonal projection onto $\Gamma_{A}$ is
@@
$\mathbf{w}_{A}=\tilde{\mathbf{w}}_{a}=\sqrt{1-a^{*}a}\oplus a,\quad\quad\mathbf{p}_{A}=\tilde{\mathbf{p}}_{a}=\tilde{\mathbf{w}}_{a}\tilde{\mathbf{w}}_{a}^{*}=\begin{pmatrix}1-a^{*}a&\sqrt{1-a^{*}a}\,a^{*}\\ a\sqrt{1-a^{*}a}&aa^{*}\end{pmatrix},$
so $\mathbf{p}$ factors through a continuous map $\tilde{\mathbf{p}}\colon\mathcal{D}(H,H^{\prime})\to\mathcal{P}(H\oplus H^{\prime})$. Restricting $\tilde{\mathbf{p}}$ to the subspace ${{}^{r}\mathcal{R}}$ of $\mathcal{D}$, we see that the identity map ${{}^{r}\mathcal{R}}\to{{}^{g}\mathcal{R}}$ is included into the following commutative square and thus is continuous:
[…]
@@
Graph-to-strong* continuity of the bounded transform.   By the definition of Riesz topology, the bounded transform $\bm{\chi}\colon\mathcal{R}(H,H^{\prime})\to\mathcal{D}(H,H^{\prime})$ is Riesz-to-norm continuous. In the proof of Theorem A we also need another type of continuity of $\bm{\chi}$.
Recall that a family of bounded operators $a_{x}\in\mathcal{B}(H,H^{\prime})$ is said to be continuous in the strong operator topology (or strongly continuous) if the family $a_{x}\xi$ of vectors in $H^{\prime}$ is continuous for every $\xi\in H$. A family $a_{x}$ is said to be continuous in the strong* operator topology if both families $a_{x}$ and $a_{x}^{*}$ are strongly continuous.
@@
The projection $\mathbf{p}_{A}$ is given by the formula
(2.6) $\mathbf{p}_{A}=\tilde{\mathbf{p}}_{a}=\begin{pmatrix}1-a^{*}a&\sqrt{1-a^{*}a}\,a^{*}\\ a\sqrt{1-a^{*}a}&aa^{*}\end{pmatrix}=\begin{pmatrix}q^{2}&qa^{*}\\ aq&aa^{*}\end{pmatrix}=\begin{pmatrix}q^{2}&b^{*}\\ b&aa^{*}\end{pmatrix}.$
@@
Choose $\xi\in H$ and $A_{0}\in\mathcal{R}(H,H^{\prime})$. Let $a_{0}$, $q_{0}$, $b_{0}$ be determined by $A_{0}$ as in (2.5). We need to show that the map ${{}^{g}\mathcal{R}}(H,H^{\prime})\to H^{\prime}$, $A\mapsto\bm{\chi}(A)\xi$ is continuous at $A=A_{0}$. Let $\varepsilon>0$. The operator $q_{0}$ is self-adjoint and injective and thus has dense range, so one can find $\eta\in H$ such that $\left\|q_{0}\eta-\xi\right\|<\varepsilon$. Let $V$ be a neighborhood of $A_{0}$ in ${{}^{g}\mathcal{R}}(H,H^{\prime})$ such that $\left\|(q-q_{0})\eta\right\|<\varepsilon$ and $\left\|(b-b_{0})\eta\right\|<\varepsilon$ for every $A\in V$. Then
$\left\|a\xi-a_{0}\xi\right\|\;<\;\left\|\left(a-a_{0}\right)q_{0}\eta\right\|+2\varepsilon\;\leqslant\;\left\|a\left(q_{0}-q\right)\eta\right\|+\left\|\left(aq-a_{0}q_{0}\right)\eta\right\|+2\varepsilon$
$\leqslant\;\left\|\left(q-q_{0}\right)\eta\right\|+\left\|\left(b-b_{0}\right)\eta\right\|+2\varepsilon<4\varepsilon$

### v3

@@ 1 Introduction
Second local obstruction.  When all the operators $\mathcal{A}_{x}$ are contained in a fixed connected component of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}$, there is another kind of local obstruction which we now describe. For a fixed $x\in X$, choose $\lambda\in\mathbb{R}$ in the resolvent set of $\mathcal{A}_{x}$. Then $\lambda$ also lies in the resolvent set of $\mathcal{A}_{y}$ for $y$ close to $x$, and the spectral projections $p_{y}=\operatorname{\mathds{1}}_{[\lambda,+\infty)}(\mathcal{A}_{y})$ determine a continuous map (a bundle projection) $p\colon\left.\mathcal{H}\right|_{Y}\to\left.\mathcal{H}\right|_{Y}$ over some neighborhood $Y$ of $x$ (here $\operatorname{\mathds{1}}_{S}$ denotes the characteristic function of a subset $S\subset\mathbb{R}$). See Proposition 7.1. A trivialization of $\mathcal{H}$ over $Y$ is adapted to $\mathcal{A}$ if and only if it is adapted to $p$ (Proposition 7.2).
@@
Suppose first that all the operators $\mathcal{A}_{x}$ are essentially negative. On the space of projections of a fixed finite rank the norm and strong topologies coincide [DD, Lemma 5], so there is an obstruction to existence of local adapted trivializations near $x$ if the rank of $p_{y}$ is not locally constant near $x$. In other words, there are eigenvalues of $\mathcal{A}_{y}$ going to $\infty$ when $y\to x$. If this is not the case and the operators $\mathcal{A}_{y}$ are uniformly bounded from above in a neighborhood of each point $x\in X$, then $\mathcal{A}$ is Riesz continuous in every trivialization of $\mathcal{H}$. The situation with essentially positive operators is completely the same, but $\infty$ and bounds from above should be replaced by -$\infty$ and bounds from below. A graph continuous family of essentially positive (resp, essentially negative) self-adjoint operators with compact resolvents may have no local adapted trivializations in a neighborhood of $x\in X$, even when the base space $X$ is an interval. See Rellich’s example [Kat, Example V-4.14], or [P3, Example 8.2] for more details.
Suppose now that all $\mathcal{A}_{x}$ are contained in the third connected component ${{}^{r}\mathcal{R}}_{K}^{\star}(\mathcal{H}_{x})$ of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}(\mathcal{H}_{x})$, which consists of operators that are neither essentially positive nor essentially negative. In general, such a family, as well as its spectral sections, may have no local adapted trivializations. See Example 7.5. However, Theorem 5 implies the following result.
Theorem 7.
@@
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$ of locally finite covering dimension. Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}_{x}$, with compact resolvents, which are neither essentially positive nor essentially negative. Then $\mathcal{A}$ admits a global adapted trivialization.
@@
Since the group ${{}^{s}{\mathcal{U}}}(H)$ is contractible, Theorem 7 implies that the identity map ${{}^{r}\mathcal{R}}^{\star}_{K}\to{{}^{g}\mathcal{R}}^{\star}_{K}$ is a weak homotopy equivalence. The stronger result, of a homotopy equivalence of this map, is proven in the previous paper of the author by a different method. See [P3], Theorems B and C.
@@
Example 7.5 below shows that Theorem 7 does not hold for general base spaces, even compact, without additional assumptions.
The author does not know whether an analogue of Theorem 7 (probably with stronger assumptions on $X$ and for maps $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}^{\star}_{F}$) holds for Fredholm operators. See Remark 7.4.
$K^{1}$ index for graph continuous families.  Let $\mathcal{A}$ be a family satisfying assumptions of Theorem 7. Then $\mathcal{A}$ admits an adapted trivialization $\tau\colon\mathcal{H}\to H_{X}$, which leads to a norm continuous map $\alpha=\bm{\chi}(\tau_{*}\mathcal{A})\colon X\to\mathcal{B}_{F}^{\star}(H)$. The homotopy class of $\alpha$ does not depend on the choice of $\tau$. This allows to define the index of $\mathcal{A}$ as the classical index of $\alpha$, $\operatorname{ind}\mathcal{A}:=\operatorname{ind}\alpha=[\alpha]\in[X,\mathcal{B}_{F}^{\star}]=K^{1}(X)$.
Another approach to the index of graph continuous families of self-adjoint Fredholm operators is provided by the previous paper [P3] of the author or by results of Joachim [J]. An approach of Ivanov [I1, I2] works even under much weaker continuity assumptions than graph continuity. The same reasoning as for the $K^{0}$ index (Section 4) shows that, for a family admitting an adapted trivialization, the index given by the formula above coincides with Ivanov’s index.
@@
Applications to elliptic operators.  The author arrived at Theorem 6 in an attempt to understand a step in the proof of a result of Melrose and Piazza. See [MP1], Proposition 1. They deal with a family $\mathcal{D}$ of self-adjoint differential operators $\mathcal{D}_{x}$ of order $1$ parametrized by $x\in X$ such that $\mathcal{D}_{x}$ acts on the fibers of a vector bundle $\mathcal{E}_{x}$ over a smooth closed manifold $\mathcal{M}_{x}$. More precisely, $\mathcal{M}\to X$ is a smooth fibration and $\mathcal{E}$ is a smooth vector bundle over $\mathcal{M}$. Melrose and Piazza showed that if the $K^{1}(X)$-index of $\mathcal{D}$ vanishes, then $\mathcal{D}$ admits a spectral section. In their proof, they claim that the $L^{2}$-spaces of the fibers, $\mathcal{H}_{x}=L^{2}(\mathcal{M}_{x};\mathcal{E}_{x})$, can be identified so that all operators act on a fixed Hilbert space $H$. Unfortunately, they did not explain how such an identification can be obtained. The author was not able to find a simple explanation of this step. The structure group of the corresponding Hilbert bundle $\mathcal{H}$ has the topology coarser than the norm topology, so Kuiper’s theorem cannot be used to trivialize $\mathcal{H}$. Dixmier-Douady theorem, on the other hand, can be used, but trivializations it provides do not take $\mathcal{D}$ to a Riesz continuous family of operators. At the same time, the proof of Melrose and Piazza uses essentially Riesz continuity of $\mathcal{D}$.
Theorem 6, as well as recent results [I2, Theorems 4.5 or 4.6] of Ivanov, allows to overcome this difficulty by turning $\mathcal{D}$ to a Riesz continuous family (though at the cost of losing the differential and geometric structure). Similarly, Theorem 3 allows to pass to a fixed Hilbert space in the proof of [MP2, Proposition 2], an odd analog of [MP1, Proposition 1].
Another framework where our results may be useful is that of elliptic boundary value problems. A continuous, in an appropriate sense, family of elliptic operators of positive order on a compact manifold with boundary and with elliptic boundary conditions leads to a graph continuous family of regular operators with compact resolvents acting between the $L^{2}$-spaces of the corresponding vector bundles. See [P1, Appendix A.5] or [BZ, Corollary A.6.3]. A fixed manifold with boundary may be replaced by a family of such manifolds, as above. Theorems 3 and 7 show that such a family of unbounded operators may be turned into a Riesz continuous family by appropriate global trivializations of the corresponding Hilbert bundles, both in the non-self-adjoint and self-adjoint case.
Acknowledgments.  This work was partially supported by ISF grants no. 431/20 and 876/20 and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant no. 101001677). It was done during my postdoctoral fellowship at the Technion – Israel Institute of Technology; the revised version was prepared during my postdoctoral fellowship at the University of Haifa.
I am grateful to N. V. Ivanov for his interest in my work and, in particular, for attracting my attention to the work of Dixmier and Douady [DD]. The exposition in this paper benefitted a lot from numerous suggestions of the anonymous referee, to whom I am deeply grateful.
@@ 2 Preliminaries
Grassmanian.  The set $\mathcal{G}r(H)$ of closed subspaces of $H$ is called the Grassmanian. An orthogonal projection may be identified with its range, which provides a bijection between $\mathcal{G}r(H)$ and the space of projections $\mathcal{P}(H)$. The Grassmanian is equipped with the topology (and even metric) induced by this bijection from the norm on $\mathcal{P}(H)$. We will use closed subspaces and projections interchangeably, depending on what is more convenient at the moment.
Canonical decomposition.  Recall a standard construction that will be used throughout the paper. The trivial Hilbert bundle over $\mathcal{G}r(H)$ with the fiber $H$ is canonically decomposed into the direct sum
$\mathcal{G}r(H)\times H=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$
of two Hilbert bundles, whose fibers over $L\in\mathcal{G}r$ are $\mathcal{H}^{+}_{L}=L$ and $\mathcal{H}^{-}_{L}=L^{\bot}$, the orthogonal complement of $L$ in $H$. This decomposition is locally trivial in the following sense: for every $L_{0}\in\mathcal{G}r(H)$ there is a continuous map $g\colon B\to\mathcal{U}(H)$, where $B$ is the open ball of radius $1$ around $L_{0}$, such that $L=g_{L}L_{0}$ for every $L\in B$. In terms of projections, this is written as $p=g_{p}p_{0}g_{p}^{*}$ for $\lVert p-p_{0}\rVert<1$, where $p_{0}$ is the projection onto $L_{0}$. See the proof of [Kat, Theorem I.6.32] for the construction of such a map $g$.
An important subspace of $\mathcal{G}r(H)$ is
$\mathcal{G}r^{\star}(H)=\left\{L\in\mathcal{G}r(H)\mid\dim L=\infty=\operatorname{codim}L\right\}.$
Let $L_{0}\in\mathcal{G}r^{\star}(H)$ be a fixed subspace. The map $\mathcal{U}(H)\to\mathcal{G}r^{\star}(H)$, $u\mapsto uL_{0}$ is a locally trivial principal bundle over a paracompact base space, whose structure group $\mathcal{U}(L_{0})\times\mathcal{U}(L_{0}^{\bot})$ is contractible. Therefore, $\mathcal{U}(H)\to\mathcal{G}r^{\star}(H)$ has a section
(2.1) $g\colon\mathcal{G}r^{\star}(H)\to\mathcal{U}(H),\;\;\text{so that}\;\;L=g_{L}L_{0}\;\;\text{for all}\;\;L\in\mathcal{G}r^{\star}(H).$
It follows that the restrictions of $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ to $\mathcal{G}r^{\star}$ are trivial Hilbert bundles with the structure group $\mathcal{U}$. In addition, contractibility of the total space $\mathcal{U}(H)$ and the fiber $\mathcal{U}\times\mathcal{U}$ implies contractibility of the base space $\mathcal{G}r^{\star}$. See [AS, proof of Lemma 3.6].
Bounded transform.  Let us recall the definition and main properties of the bounded transform. See, e.g., [S], Theorem 7.5 for the detail.
@@
defines the injective map from the set $\mathcal{R}(H,H^{\prime})$ of regular operators to the closed unit ball
$\mathcal{D}(H,H^{\prime})=\left\{a\in\mathcal{B}(H,H^{\prime})\mid\lVert a\rVert\leqslant 1\right\}.$
@@
Let $\mathbf{p}_{A}$ denote the orthogonal projection onto $\Gamma_{A}$,
@@
(2.5) $\mathbf{w}_{A}=\sqrt{1-a^{*}a}\oplus a,\quad\quad\mathbf{p}_{A}=\begin{pmatrix}1-a^{*}a&\sqrt{1-a^{*}a}\,a^{*}\\ a\sqrt{1-a^{*}a}&aa^{*}\end{pmatrix}$,
so $\mathbf{p}$ factors through a continuous map $\mathcal{D}(H,H^{\prime})\to\mathcal{P}(H\oplus H^{\prime})$. Restricting this map to the subspace $^{r}\mathcal{R}$} of $\mathcal{D}$, we see that the identity map ${{}^{r}\mathcal{R}}\to{{}^{g}\mathcal{R}}$ is continuous.
@@
Graph-to-strong* continuity of the bounded transform.  By the definition of Riesz topology, the bounded transform $\bm{\chi}\colon\mathcal{R}(H,H^{\prime})\to\mathcal{D}(H,H^{\prime})$ is Riesz-to-norm continuous. In the proof of Theorem 1 we also need another type of continuity of $\bm{\chi}$.
Let ${{}^{s}{\mathcal{B}}}(H,H^{\prime})$ denote the space of bounded operators $H\to H^{\prime}$ equipped with the strong operator topology. Recall that this topology is defined as the weakest topology for which the map ${{}^{s}{\mathcal{B}}}(H,H^{\prime})\to H^{\prime}$, $a\mapsto a\xi$ is continuous for every $\xi\in H$. The strong* operator topology on the same set is the weakest topology for which both $a\mapsto a\xi$ and $a\mapsto a^{*}\xi$ are continuous for every $\xi\in H$.
@@
In these terms, the second equality of (2.5) is written as
(2.7) $\mathbf{p}_{A}=\begin{pmatrix}q^{2}&qa^{*}\\ aq&aa^{*}\end{pmatrix}=\begin{pmatrix}q^{2}&b^{*}\\ b&aa^{*}\end{pmatrix}$.
@@
Choose $\xi\in H$ and $A_{0}\in\mathcal{R}(H,H^{\prime})$. Let $a_{0}$, $q_{0}$, $b_{0}$ be determined by $A_{0}$ as in (2.6). We need to show that the map ${{}^{g}\mathcal{R}}(H,H^{\prime})\to H^{\prime}$, $A\mapsto\bm{\chi}(A)\xi$ is continuous at $A=A_{0}$. Let $\varepsilon>0$. The operator $q_{0}$ is self-adjoint and injective and thus has dense range, so one can find $\eta\in H$ such that $\lVert q_{0}\eta-\xi\rVert<\varepsilon$. Let $V$ be a neighborhood of $A_{0}$ in ${{}^{g}\mathcal{R}}(H,H^{\prime})$ such that $\lVert(q-q_{0})\eta\rVert<\varepsilon$ and $\lVert(b-b_{0})\eta\rVert<\varepsilon$ for every $A\in V$. Then
$\lVert a\xi-a_{0}\xi\rVert\;<\;\lVert\left(a-a_{0}\right)q_{0}\eta\rVert+2\varepsilon\;\leqslant\;\lVert a\left(q_{0}-q\right)\eta\rVert+\lVert\left(aq-a_{0}q_{0}\right)\eta\rVert+2\varepsilon$
$\leqslant\;\lVert\left(q-q_{0}\right)\eta\rVert+\lVert\left(b-b_{0}\right)\eta\rVert+2\varepsilon<4\varepsilon$